Suppose a car rental firm wants to estimate the average number of kilometers traveled per day by each of its cars rented in a certain city. A random sample of 20 cars rented in that city reveals that the sample mean travel distance per day is 85.5 kilometers, with a population standard deviation of 19.3 kilometers. Compute a 99% confidence interval to estimate Q. (2 points) Interpret your answer. (1 point)​


1
Expert's answer
2022-06-06T00:16:27-0400

The critical value for "\alpha = 0.01, df=n-1=19" degrees of freedom is "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=2.860935."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(85.5- 2.860935\times\dfrac{19.3}{\sqrt{20}},""85.5+ 2.860935\times\dfrac{19.3}{\sqrt{20}})"




"=(73.153, 97.847)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "73.153 < \mu < 97.847," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(73.153, 97.847)."